OpenStudy (anonymous):

What is the vertex of the graph of the function below? y = x2 - 8x + 12 A. (2, 0) B. (4, -4) C. (4, 0) D. (2, -4)

OpenStudy (imstuck):

Start like this so you can begin to complete the square on this:

OpenStudy (anonymous):

Use the vertex formula: X=-b/(2a) Y=f(above value)

OpenStudy (imstuck):

\[y-12=x ^{2}-8x\]Now complete the square on the x terms like this:

OpenStudy (igreen):

\(\dfrac{-b}{2a}\)

OpenStudy (igreen):

In this case: a = 1 b = -8

OpenStudy (imstuck):

\[y-12=x ^{2}-8x+16\]and of course you have to add the 16 to the other side too, to get this:

OpenStudy (igreen):

\(\dfrac{8}{2(1)}\) \(\dfrac{8}{2}\) \(4\)

OpenStudy (imstuck):

\[y-12+16=(x-4)^{2}\]\[y+4=(x-4)^{2}\]

OpenStudy (igreen):

x-intercept of your vertex is 4.

OpenStudy (imstuck):

Looking at that equation, you can see that the x coordinate is +4, and the y coordinate is -4. So your vertex is (4, -4). B from your choices.

OpenStudy (igreen):

Plug in 4 for x in the equation: y = x^2 - 8x + 12 y = (4)^2 - 8(4) + 12 y = 16 - 32 + 12 y = -16 + 12 y = -4

OpenStudy (igreen):

So your vertex is (4, -4)

OpenStudy (imstuck):

If you learn how to complete the square for these, you can find everything you need to know about them. The directrix, the focus, the vertex, everything!

OpenStudy (igreen):

You can give medals by clicking 'Best Response'. @mikerinks

OpenStudy (imstuck):

hint, hint... ; )